# The anomalous magnetic moment of the muon: status and perspectives

### v1

@@ 5.1.1 $e^{+}e^{-}\to\pi^{+}\pi^{-}(\gamma)$
An important aspect of resolving the puzzling situation in $e^{+}e^{-}\to\pi^{+}\pi^{-}$, see Fig. 5(A), involves new data and analyses. In the meantime, such new results have become available from BaBar, based on data that had never been analyzed before, and SND. Preliminary results suggest that the new BaBar analysis almost coincides with their previous result, while the new results from SND move closer to CMD-3 compared to their earlier measurement [146]. Similarly to BaBar, a large amount of data from the KLOE experiment had never been analyzed, and an effort to do so is ongoing. Finally, new data are expected from BESIII and Belle II.
In all cases, special emphasis is put on the role of radiative corrections and MC generators [161, 162, 163, 164, 165, 166, 167, 168], to scrutinize whether relevant effects were missed in previous analyses and ensure state-of-the-art input for ongoing ones. In fact, one of the cross checks performed by CMD-3 [147, 148] concerns the forward–backward asymmetry in $e^{+}e^{-}\to\pi^{+}\pi^{-}$, which is $C$-odd, thus canceling in the total cross section, but still constitutes an interesting test case. It was observed that the MC simulation based on scalar QED (sQED) for the pion states could not reproduce the data, a mismatch that could be resolved by structure-dependent radiative corrections [162], as shown in Fig. 6(A). In particular, the first diagram contributes to the asymmetry, interfering with the tree-level amplitude, and is only poorly represented by the point-like, sQED approximation. In the language of dispersion relations [163], this topology involves the general pion Compton tensor [171, 172, 173, 174, 175, 176], but the leading contribution originates from the pion pole. In the end, it is the interplay of pole and dispersive corrections, as a remnant of the cancellation of the IR singularity, that displays an enhancement with the $\rho(770)$ resonance, see Fig. 6(B). This resonance effect causes a sizable enhancement beyond the point-like limit, explaining the observation by CMD-3.
@@ 5.1.2 Hadronic $\tau$ decays
In Ref. [22] three main effects were identified that cast doubts on the reliability of the currently available isospin-breaking corrections. First, the matching of short-distance corrections with hadronic matrix elements involved an ${\mathcal{O}}(\alpha/\pi)$ scheme ambiguity [178], which, however, is being addressed using input from lattice QCD for the nonperturbative matching [179, 180]. Second, resonance-enhanced virtual corrections could become important, see second diagram in Fig. 6(A), and indeed these effects were evaluated recently [169, 170], again leading to sizable corrections beyond the point-like limit, see Fig. 6(C). Third, isospin breaking in the matrix elements needs to be quantified, which constitutes the most challenging task, but could be addressed using lattice QCD [181] and dispersive techniques [182]. In addition to these crucial theory advances, also new data especially on the $\tau\to\pi\pi\nu_{\tau}$ spectrum would be extremely valuable, as could become available at Belle II [183].

### v2

@@ 5.1.1 $e^{+}e^{-}\to\pi^{+}\pi^{-}(\gamma)$
An important aspect of resolving the puzzling situation in $e^{+}e^{-}\to\pi^{+}\pi^{-}$, see Fig. 5(A), involves new data and analyses. In the meantime, such new results have become available from BaBar, based on data that had never been analyzed before, and SND. Preliminary results [173] suggest that the new BaBar analysis almost coincides with their previous result, while the new results from SND move closer to CMD-3 compared to their earlier measurement [147]. Similarly to BaBar, a large amount of data from the KLOE experiment had never been analyzed, and an effort to do so is ongoing. Finally, new data are expected from BESIII and Belle II.
In all cases, special emphasis is put on the role of radiative corrections and MC generators [162, 163, 164, 165, 166, 167, 168, 169, 170], to scrutinize whether relevant effects were missed in previous analyses and ensure state-of-the-art input for ongoing ones. In fact, one of the cross checks performed by CMD-3 [148, 149] concerns the forward–backward asymmetry in $e^{+}e^{-}\to\pi^{+}\pi^{-}$, which is $C$-odd, thus canceling in the total cross section, but still constitutes an interesting test case that could have implications for $C$-even contributions in measurements using initial-state radiation, see Fig. 6(A). It was observed that the MC simulation based on scalar QED (sQED) for the pion states could not reproduce the data, a mismatch that could be resolved by structure-dependent radiative corrections [163], as shown in Fig. 6(A). In particular, the first diagram contributes to the asymmetry, interfering with the tree-level amplitude, and is only poorly represented by the point-like, sQED approximation. In the language of dispersion relations [164], this topology involves the general pion Compton tensor [174, 175, 176, 177, 178, 179], but the leading contribution originates from the pion pole. In the end, it is the interplay of pole and dispersive corrections, as a remnant of the cancellation of the IR singularity, that displays an enhancement with the $\rho(770)$ resonance, see Fig. 6(B). This resonance effect causes a sizable enhancement beyond the point-like limit, explaining the observation by CMD-3.
@@ 5.1.2 Hadronic $\tau$ decays
In Ref. [22] three main effects were identified that cast doubts on the reliability of the currently available isospin-breaking corrections. First, the matching of short-distance corrections with hadronic matrix elements involved an ${\mathcal{O}}(\alpha/\pi)$ scheme ambiguity [181], which has now been addressed [182] using input from lattice QCD for the nonperturbative matching [183, 184]. Second, resonance-enhanced virtual corrections could become important, see second diagram in Fig. 6(A), and indeed these effects were evaluated recently [171, 172], again leading to sizable corrections beyond the point-like limit, see Fig. 6(C). Third, isospin breaking in the matrix elements needs to be quantified, which constitutes the most challenging task, but could be addressed using lattice QCD [185] and dispersive techniques [186]. In addition to these crucial theory advances, also new data especially on the $\tau\to\pi\pi\nu_{\tau}$ spectrum would be extremely valuable, as could become available at Belle II [187].